Vector A has magnitude 4.0 units and vector B has magnitude 5.0 units. The angle between A and B is 60.0°. What is the magnitude of A+B ?


1
Expert's answer
2021-10-20T10:41:40-0400

The magnitude is


"|A+B|=\sqrt{(A+B\cos\theta)^2+(B\sin\theta)^2},\\
A+B|=7.8\text{ units}."